A Complete Bibliography of Publications in Journal of Fish Diseases (1970–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

19 September 2022 Version 1.03

Title word cross-reference

14 [KKS^{+}79]. _{2} [ESB79, SKH^{+}79].

-like [LJ78].

2 [DCT78].

7138 [MGP78].

aberrant [HSR78]. Acanthocephala [Ami78, Val79]. Acholeplasma [ENB79]. acid [KKS^{+}79, KKT^{+}79]. action [GGB79]. activity [LS79, OCF78]. aculeatus [DL78]. adult [BBN79, Mil79]. Aeromonas [BCC78, HSR78, HEH79, RCC78]. aetiology [BCC^{+}79]. affected [RH78]. after [KK78, WQCO79]. against [EA79b, EA79a, Mca78]. age [RG78]. agent [Mey79]. agents [BR79]. agglutination [OCF78]. airbreathing [MAJK79]. Alaskan [MGMW79]. alevins [Arm79]. Algal [BT79, Edw78]. alpinus [RP78]. American [HKMS79]. among [HSR78]. amongst [FB79]. anaemia [Smi79]. analysis [Wil78]. Anguilla [PP79]. anguillarum [EA78]. anomala [DL78]. anomalies [PBH79]. antibiotic [KKT^{+}79]. antibodies [DT79, IH79]. antibody [Mca78]. antigen [ARD79]. antigenic [DCT78]. antiserum [OCF78]. Apatemon [RG78]. Aphanopus [Agi78]. areas [BSS78]. Aroclor [LS79]. aspects [FB79]. assay [ARD79]. associated [BSHR79]. association [Buc78]. asymmetry [RGC79]. atherinodi [HW78]. atherinoides [HW78]. Atlantic [Agi78, Dun78, EW78, PPS^{+}78, RHH78, RH79, SPKM79]. Attachment [Mil79, FSH78]. aumtus [KKS^{+}79]. aurata [PSC78]. auratus [CS79, DL79, RCC78]. Australia [HWL79]. Ayres [BT79].

baccatus [Som78]. bacteria [KK78]. Bacterial [RH78, BSHR79]. bacterium [BSS78]. Balantidium [MR78]. bass [HEH79]. Bath [EA79b, Wal79]. Bath-immunization [EA79b]. Bay [HWL79, Mey79, Val79]. behaviour [BC79]. between [Rob79]. Bilateral [RGC79]. Billet [MAJK79]. biochemical [MAJK79]. biological [DCT78]. biotransformation [KKS^{+}79]. bivalve [Buc78]. Bloch [MAJK79]. Book [Ano78a, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano78b]. borelli [DL79]. Boschma [BC79, PC78]. Bothnia [Val79]. Bothriocephalus [SG79]. Brachydanio [Mil79]. brown [FM78, RP78, RP79, CNR79]. Buccinum [Som78]. Buchanan [Mil79].

C [KKS^{+}79]. C-piromidic [KKS^{+}79]. cancer [HWL79]. captive [BFMS79]. Carassius [CS79, KKS^{+}79, RCC78]. Carassiw [DL79]. carbo [Agi78]. carp [Ahn78, BSHR79]. carpio [Ahn78]. caryophyllid [Hay79]. case [CS79]. catfish [BCC78, MGP78]. Catostomus [Hay79]. cause [BW79]. caused [Hay79, SG79]. cell [ENB79, KK79, PSC78]. cells [DCT78, HLF78]. Cellular [ARD79]. centres [Agi79]. cephalus [Arm79]. Ceratomyxa [YS79]. ceroidosis [Smi79]. Cestoda [SG79]. Changes [BCC78, RP79, DL79, MAJK79]. Channa [MAJK79]. channel [BCC78, MGP78]. char [RP78]. Chen [MR78]. Chilodonella [HKMS79]. chlamydial [Mey79]. chronic [RH78]. Ciliata [HKMS79]. Cirripedia [BC79, PC78]. clams [Mey79]. clinical [RRB79]. closed [CSE79]. clupeaformis [RGC79]. CO [ESB79, SKH^{+}79]. coastal [Arm79]. Coho [HHS79, OCF78]. coli [RCC78]. collected [SPKM79]. commercial [BC79, PC78]. commersoni [Hay79]. comparison [BSS78]. Complement [DT79, IH79]. complement-fixation [IH79]. computer [RGC79]. concentrations [SKH^{+}79]. content [ESB79]. Coregonm [RGC79, Val79]. Costa [Buc78]. Costiasis [EW78]. crab [BC79, PC78]. critical [Wil78]. Cryptocaryon [CNR79]. cryptocaryoniasis [CNR79]. Ctenopharyngodon [MR78]. ctenopharyngodonis [MR78]. Culaea [RG78]. cultivated [MW79]. culture [KK79]. cultured [KK79, Kir78, KK78, RB79]. cultures [ENB79]. cycle [CNR79]. cyprinid [SG79]. Cyprinus [Ahn78]. Cytological [Buc78].

danilewskyi [DL79]. deep [Agi78]. deep-sea [Agi78]. degeneration [Smi79]. demersal [HWL79]. dependent [DT79]. dermal [LJ78]. desiccation [WQCO79]. detection [IH79]. developmental [PBH79]. Dicauda [HW78]. dietary [LS79]. diets [Smi79]. differences [PPS^{+}78]. different [BSS78, NTJH79]. digenean [MAJK79, Mil79]. digestive [Buc78, Mey79, Som78]. Diplostomum [RGC79]. disease [Ami79, FN78, FB79, HEH79, LJ78]. diseased [Agi79]. diseases [Arm79, MGMW79]. Distribution [KKS^{+}79, RG78, RGC79, RP78]. diverticular [Mey79]. Dover [MW79]. due [HSR78]. during [HWL79]. dysplasia [CWHG79].

East [Agi78]. Echinorhynchidae [Ami78]. Echinorhynchus [Ami78]. Ecological [PC78]. ectocommensal [FSH78]. ectoparasitic [Mil79]. eel [PP79]. eels [Mca78]. Effect [Ami78, HFK79, LS79, MAJK79, CNR79, KK79, OCF78, SKH^{+}79]. Effects [ENB79]. egg [Wal79]. eggs [FG78]. Egtved [DT79]. Electron [LJ78, Fer79, YS79]. elevated [SKH^{+}79]. emphasis [FSH78]. endotoxin [RCC78]. England [BFMS79]. Enteric [BSS78]. Epidemiological [FB79]. epidermal [PP79]. epitheliocystis [PSC78]. epithelium [Mey79, PSC78]. epizootic [BCC^{+}79]. epizootiology [Wya78]. erectus [Bla79]. Errata [Ano79f]. Escherichia [RCC78]. estuary [SPKM79]. European [PP79]. evaluation [KKT^{+}79]. Evidence [Dun78]. excretion [KKS^{+}79]. Exophiala [BW79, RHH78]. Exophiala-like [BW79]. experimental [KK79]. experimentally [HEH79]. exposed [CWHG79]. exposure [BR79, SKH^{+}79]. eyed [FG78]. eyes [RGC79].

farmed [EA79b, RH78, Rob79]. fate [KK78]. features [DCT78]. Febrile [RCC78]. fed [Smi79]. feeding [Mil79]. fibrosarcoma [Dun78]. Fin [MW79]. Fine [FSH78]. Finland [LJ78]. fish [Agi78, Agi79, BW79, BFMS79, BR79, CNR79, CWHG79, Edw78, ENB79, HWL79, HLF78, HKMS79, KKT^{+}79, MAJK79, Mil79, PSC78, Wil78]. fishes [Ami79, BO79, MGMW79, SG79]. fixation [IH79]. Flatfish [PPS^{+}78]. fluviatilis [And79, BCC^{+}79]. following [BSHR79]. Foreword [Sni78]. formalin [GGB79]. found [Buc78, HWL79]. free [KK79]. Frequency [RP78, RG78]. fresh [HR79]. fry [Fer79]. fungus [Agi78, BW79]. Furanace [EA79a, MGP78]. Further [BCC^{+}79]. furunculosis [FM78].

gairdneri [ARD79, CSE79, ESB79, EA78, EA79b, EA79a, FN78, FB79, Fer79, FG78, HR79, HFK79, IH79, LS79, RH78, RB79, RRB79, SKH^{+}79, Smi79]. gaivdneri [Wal79]. Gasterosteus [DL78]. geographic [BSS78]. gill [PSC78]. gland [Buc78, Som78]. Glugea [Bla79, DL78]. Glugeidae [Bla79]. goitre [CS79]. goldfish [CS79, DL79, KKS^{+}79, RCC78]. gowkongensis [SG79]. gracilis [RG78]. granifera [BC79, PC78]. grasscarp [MR78]. Great [Mey79]. green [GGB79]. growth [PP79]. Gulf [BO79].

haematological [MAJK79]. haematopoietic [CSE79, HFK79]. Haematozoa [BO79]. Haemogregavina [Kir78]. haemolytic [ARD79]. Hamilton [Mil79]. Hamilton-Buchanan [Mil79]. handling [BSHR79]. hard [Mey79]. hatchery [RP78]. hatchery-reared [RP78]. Henneguy [Rob79]. Henneguya [And79]. hepatoma [Wal79]. Hepatomas [SPKM79]. heraldi [Bla79]. herbicide [CWHG79]. Herpesvirus [BM78, RB78]. Hexamita [Fer79]. hexasticha [HKMS79]. high [ESB79]. Hippocampus [Bla79]. Histopathological [DL79, RRB79, RB78]. Histopathology [FM78, HWL79, HEH79, HR79, Som78]. homogeneous [RG78]. Host [EA78, Rob79, Ami78, NTJH79, PSC78]. Host-parasite [Rob79]. Host-specific [EA78]. Hudson [SPKM79]. Humoral [Mca78]. hydrophila [BCC78, HEH79, RCC78]. Hypophthalmichthys [BSHR79]. hypselobagri [MAJK79].

Ichthyophonus [Agi78]. Ichthyophonus-like [Agi78]. Ichtyobodo [Rob79]. Ictalurus [BCC78, MGP78]. idella [MR78]. immune [ARD79]. immunization [EA79b]. inconstans [RG78]. Induction [BR79, Wal79]. indueed [BCC78]. infected [HEH79, MR78, RP79]. Infection [Agi78, BCC78, BSHR79, BT79, BM78, DL78, HFK79, KK79, KK78, RHH78, RP78, RB78]. infections [DL79, Edw78, HLF78, RGC79]. Infectious [BFMS79, CSE79, DCT78, FG78, HFK79, HLF78, NTJH79]. infectivity [WQCO79]. influence [PP79]. inoculation [KK78]. insidiosus [Wya78]. interactions [Rob79]. Intestinal [MR78, Hay79, Mca78]. Ionic [ESB79]. IPN [BFMS79]. IPNV [HLF78, NTJH79]. Ireland [FB79]. iron [Agi79]. irritans [CNR79]. isolated [EA78]. isolates [BSS78, NTJH79]. isolation [FG78]. Isoparorchis [MAJK79].

kept [ESB79]. kidney [FN78, FB79]. Kiernik [HKMS79]. Kirtland [RG78]. kisutch [HHS79, OCF78].

L. [Ahn78, And79, Arm79, BC79, BM78, BCC^{+}79, BBN79, CS79, Dun78, DL78, DL79, EA78, EW78, FM78, HSR78, KKS^{+}79, Kir78, MW79, Mey79, PSC78, PP79, PC78, PBH79, RCC78, RHH78, RP78, RB78, RP79, RH79, Som78]. Lacépède [HEH79]. Laécpède [Hay79]. lagoon [Arm79]. laidlawii [ENB79]. lake [RGC79, HSR78]. lamprey [PBH79]. largemouth [HEH79]. Laveran [DL79]. lesion [WQCO79]. lesions [BR79, MGP78, MR78, PBH79]. Light [YS79, BR79]. like [Agi78, LJ78, BW79, BR79]. lipid [LS79]. lipoid [Smi79]. Lipophoromatosis [BBN79]. liver [Smi79]. Liza [PSC78]. Llyn [And79]. localization [Ami78]. LOWQ [Agi78]. Lymphocystis [Ami79, WQCO79]. Lymphosarcoma [RH79]. lyophilization [WQCO79].

Macfarlane [Mca78]. macrophage [Agi79]. maculatus [BT79]. malachite [GGB79]. marine [BO79, BW79, Buc78, CNR79, MGMW79, RB79]. marinus [PBH79]. Mass [BSHR79, HSR78]. maturation [Ami78]. maximus [BM78, Kir78, RB78]. melano [Agi79]. melano-macrophage [Agi79]. meningitis [RH78]. Mercenaria [Mey79]. Mesnil [DL79]. metacercarial [RGC79]. method [EA79b, GGB79, IH79]. Mexico [BO79]. microbial [MGMW79]. microcytic [Smi79]. Microgadus [SPKM79]. Micropterus [HEH79]. microscopic [YS79]. microscopical [Fer79]. microscopy [LJ78]. Microsporida [Bla79]. minnow [HW78]. minnows [HSR78]. mirex [LS79]. Mitchill [RGC79]. Moche [Arm79]. mode [FSH78]. model [GGB79]. molitrix [BSHR79]. mollusc [Buc78]. Moniez [DL78]. monitored [ARD79]. Moore [Fer79]. Morbid [BC79]. mortalities [BSHR79, Fer79]. mortality [HSR78]. Mugil [Arm79]. Müller [Ami78, Val79]. mullet [Arm79]. multiplication [Ahn78]. muscle [ESB79]. mycoplasma [ENB79]. mycosis [BW79]. Myxobolus [Wya78]. myxosporida [YS79, And79, Wya78]. myxosporidan [HW78].

n [Bla79, HW78]. n.g [HW78]. n.sp [Kir78]. nasus [Val79]. naturally [HEH79]. necator [Rob79]. necrosis [BFMS79, CSE79, DCT78, FG78, HFK79, HLF78, LJ78, MW79, NTJH79, RH78]. Neoechinorhynchus [Val79]. Nephrocalcinosis [SKH^{+}79, HR79]. neutralization [DT79]. Nicoll [Som78]. Nifurpirinol [MGP78]. Noble [YS79]. non [DCT78, EA79b]. non-pathogenic [DCT78]. non-stressing [EA79b]. normal [Agi79]. North [Agi78, HKMS79]. northern [BO79, FB79]. Norway [HSR78]. Notorynchus [BT79]. Notropis [HW78].

O [ARD79]. O-antigen [ARD79]. Observations [Arm79, Hay79, BM78, Fer79, PC78, RB78, YS79]. observed [RGC79]. occurrence [And79, BFMS79]. Oncorhynchus [HHS79, OCF78]. oncovirus [Dun78]. opisthorchis [Mca78]. organisms [PSC78]. other [PBH79].

P [MGP78]. P-7138 [MGP78]. Pacific [PPS^{+}78]. Pallas [Val79]. pancreatic [BFMS79, DCT78, FG78, HLF78, NTJH79, RH78]. Pansteatitis [RRB79]. papillomas [PP79]. Papillomatosis [RB79]. parameters [RP79]. parasite [PC78, Rob79]. parasites [Arm79]. parasiticides [GGB79]. parasitized [BC79]. particular [MAJK79]. passive [ARD79]. pathogenesis [HHS79, KK79, KK78]. pathogenic [DCT78]. pathogenicity [EA78]. Pathology [SG79, HR79, Hay79]. patialense [Mil79]. patterns [RP78]. PCB [LS79]. pelagicus [BC79, PC78]. pellucidus [RG78]. penaeid [FSH78]. Perca [And79, BCC^{+}79]. perch [And79, BCC^{+}79]. peritrichous [FSH78]. Perry [Bla79]. Persistent [HLF78]. Peru [Arm79]. Petromyzon [PBH79]. phage [Buc78]. Philip [HWL79]. phototoxic [BR79]. Phoxinus [HSR78]. piromidic [KKS^{+}79, KKT^{+}79]. plaice [BBN79]. plaque [ARD79]. platessa [BBN79]. Pleuronectes [BBN79]. Pollachius [EA78]. populations [RG78]. Port [HWL79]. Portunus [BC79, PC78]. practical [EA79b]. Pratt [Wya78]. Preliminary [Mey79, Smi79]. preparations [WQCO79]. presence [BR79]. prevention [Smi79]. production [Mca78]. Proliferative [FN78, FB79]. Protozoa [HKMS79, Wya78]. Pseudophyllidea [SG79]. psorospermica [And79]. punctatus [MAJK79, MGP78]. punetatus [BCC78].

Rafinesque [BCC78, HW78, MGP78]. rainbow [ARD79, CSE79, ESB79, EA78, EA79b, EA79a, FN78, FB79, Fer79, FG78, HR79, HFK79, IH79, LS79, RH78, RB79, RRB79, SKH^{+}79, Smi79, Wal79]. ramada [PSC78]. rancid [Smi79]. range [NTJH79]. reaction [DL78]. reared [RP78]. record [Edw78]. red [HEH79]. red-sore [HEH79]. redmouth [BSS78]. reference [MAJK79]. related [Fer79, MGP78]. relation [PSC78]. replication [ENB79]. report [Smi79]. reproductive [CNR79]. rerio [Mil79]. reserves [LS79]. response [ARD79]. responses [RCC78]. review [Edw78]. Reviews [Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano78a, Ano78b]. Rhizocephala [BC79, PC78]. Richardson [ARD79, CSE79, ESB79, EA78, EA79b, EA79a, FN78, FB79, Fer79, FG78, HR79, HFK79, IH79, LS79, RH78, RB79, RRB79, SKH^{+}79, Smi79, Wal79]. rickettsia [Buc78]. Rio [Arm79]. Risso [PSC78]. River [SPKM79]. role [Agi79]. RTG [DCT78]. RTG-2 [DCT78]. ruckeri [ARD79]. rutili [Val79].

Sacculina [BC79, PC78]. sachai [Kir78]. saithe [EA78]. salar [Dun78, EW78, RHH78, RH79]. salinity [CNR79, PP79]. Salmo [ARD79, CSE79, Dun78, ESB79, EA78, EA79b, EA79a, EW78, FM78, FN78, FB79, Fer79, FG78, HR79, HFK79, IH79, LS79, RHH78, RP78, RP79, RH79, RH78, RB79, RRB79, SKH^{+}79, Smi79, Wal79]. salmoides [HEH79]. salmon [Dun78, EW78, HHS79, LJ78, OCF78, RHH78, RH79]. salmonicida [HSR78]. salmonid [BR79, HLF78, Wil78]. salmonids [DCT78, Rob79]. salmonis [Ami78, Fer79, RHH78]. salt [EA79a]. Salvelinus [RP78]. sand [BC79, PC78]. Saprolegnia [RP78, RP79]. Saprolegnia-infected [RP79]. Saprolegnias [Wil78]. scabbard [Agi78]. Scanning [Fer79]. Scophthabnus [Kir78]. scophthalmi [BM78, RB78]. Scophthalmus [BM78, RB78]. Scotland [Kir78]. sea [Agi78, EA79b, PBH79]. seahorse [Bla79]. seawater [EW78]. selected [DCT78, ENB79]. semi [CSE79]. semi-closed [CSE79]. sensu [Val79]. Seriola [KK79, KK78]. serum [RP79]. sevengill [BT79]. shark [BT79]. shasta [YS79]. shrimp [FSH78]. silver [BSHR79]. simple [WQCO79]. simulation [RGC79]. size [RG78]. skin [MW79, PPS^{+}78]. smelts [EW78]. sole [MW79]. Solea [MW79]. Soparkar [Mil79]. sore [HEH79]. South [Mey79]. sp [Bla79, FSH78, HW78, KK78]. Sparus [PSC78]. spawning [Ami78]. species [Buc78, HWL79]. specific [EA78, IH79]. spined [DL78]. spleen [BCC78]. sporogenesis [YS79]. spp [KK79, KK78]. spring [Ahn78]. Stephanochasmus [Som78]. stickleback [DL78, RG78]. storage [Agi79]. stored [WQCO79]. strain [DCT78, HSR78]. strains [EA78]. streptococcal [KK79, KK78]. Streptococcus [KK78]. stressing [EA79b]. structure [FSH78, PP79, RG78]. Studies [BM78, CNR79, HHS79, KK79, KK78, NTJH79, RB78, Wya78, Buc78, BCC^{+}79, Mey79]. study [KKT^{+}79, PSC78, RRB79]. sucker [Hay79]. superficial [HWL79]. survey [HWL79]. Svärdson [Val79]. swimbladder [Dun78]. synergistic [GGB79]. system [CSE79, GGB79]. systemic [BW79].

tapeworms [Hay79]. technique [Wal79]. Tegid [And79]. Tellina [Buc78]. Telogaster [Mca78]. temperature [CNR79, HFK79, OCF78]. tenuis [Buc78]. Teratoid [PBH79]. test [ARD79]. testing [GGB79]. tetracotyles [RG78]. their [PSC78]. Thelohan [And79]. three [DL78, Hay79]. three-spined [DL78]. thyroid [LS79]. Tissue [DL78, Edw78]. tomcod [SPKM79, SPKM79]. transmission [Fer79]. Transversotrema [Mil79]. treatment [CS79]. trematode [Mca78]. Trichiuridae [Agi78]. trifluralin [CWHG79]. trout [ARD79, CSE79, DT79, ESB79, EA78, EA79b, EA79a, FM78, FN78, FB79, Fer79, FG78, HR79, HFK79, IH79, LS79, RP78, RP79, RH78, RB79, RRB79, SKH^{+}79, Smi79, Wal79]. trutta [FM78, RP78, RP79]. Trypanoplasma [DL79]. Trypanosoma [DL79]. tumour [HWL79]. Tumours [MGMW79, PPS^{+}78]. turbot [BM78, Kir78, RB78]. Tveitevatn [HSR78].

UDN [BR79, LJ78]. UDN-like [BR79]. ulcerative [LJ78]. Ultrastructural [PSC78, BM78]. ultraviolet [BR79]. uncertain [BCC^{+}79]. undatum [Som78]. unidentified [ENB79]. Uptake [Ahn78]. urine [ESB79]. use [EA79a]. using [GGB79].

vaccinating [EA79b]. Valenciennes [BSHR79, MR78]. variant [DCT78, NTJH79]. Vertebral [CWHG79]. Vibrio [EA78, IH79]. Vibrio-specific [IH79]. vibriosis [EA79b, EA79a, HHS79]. Victoria [HWL79]. viraemia [Ahn78]. virens [EA78]. virus [Ahn78, BFMS79, DT79, DCT78, FG78, HFK79, HLF78, NTJH79, WQCO79]. viruses [ENB79]. vitro [GGB79, KKT^{+}79]. vivo [KKT^{+}79].

Walbaum [HHS79, OCF78, SPKM79]. Wales [And79, BFMS79]. warmwater [HKMS79]. water [ESB79, EA79a, HR79, HFK79]. waters [MGMW79]. white [Hay79]. whitefish [RGC79, Val79]. wild [BFMS79, RP78]. Windermere [Wil78]. Wisconsin [Ami79]. Wyatt [Wya78].

Yamaguti [RG78]. Yea [SG79]. yellowtails [KK79, KK78]. Yersinia [ARD79]. York [Mey79, SPKM79]. young [CWHG79].

Zealand [Mca78]. zebra [Mil79]. Zoothamnium [FSH78].

References

Agius:1978:IIL

[Agi78] C. Agius. Infection by an Ichthyophonus-like fungus in the deepsea scabbard fish Aphanopus carbo (LOWQ) (Trichiuridae) in the North East Atlantic. Journal of Fish Diseases, 1(2):191–193, April 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Agius:1979:RMM

[Agi79] C. Agius. The role of melano-macrophage centres in iron storage in normal and diseased fish. Journal of Fish Diseases, 2(4):337–343, July 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahne:1978:UMS

[Ahn78] W. Ahne. Uptake and multiplication of spring viraemia of carp virus in carp, Cyprinus carpio L. Journal of Fish Diseases, 1(3): 265–268, July 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amin:1978:EHS

[Ami78] Omar M. Amin. Effect of host spawning on Echinorhynchus salmonis Müller, 1784 (Acanthocephala: Echinorhynchidae) maturation and localization. Journal of Fish Diseases, 1(2):195–197, April 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amin:1979:LDW

[Ami79] O. M. Amin. Lymphocystis disease in Wisconsin fishes. Journal of Fish Diseases, 2(3):207–217, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Andrews:1979:OHP

[And79] C. Andrews. The occurrence of Henneguya psorospermica Thelohan, 1895 (Myxosporida) on perch, Perca fluviatilis L., from Llyn Tegid, Wales. Journal of Fish Diseases, 2(1):27–33, January 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1978:BRa

[Ano78a] Anonymous. Book reviews. Journal of Fish Diseases, 1(1):133– 134, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1978:BRb

[Ano78b] Anonymous. BOOK REVIEWS. Journal of Fish Diseases, 1(2): 199–203, April 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1978:BRc

[Ano78c] Anonymous. Book reviews. Journal of Fish Diseases, 1(3):275–282, July 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1978:BRd

[Ano78d] Anonymous. Book reviews. Journal of Fish Diseases, 1(4):395– 396, October 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1979:BRa

[Ano79a] Anonymous. Book reviews. Journal of Fish Diseases, 2(1):81–83, January 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1979:BRb

[Ano79b] Anonymous. Book reviews. Journal of Fish Diseases, 2(2):165–168, March 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1979:BRc

[Ano79c] Anonymous. Book reviews. Journal of Fish Diseases, 2(3):261–262, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1979:BRd

[Ano79d] Anonymous. Book reviews. Journal of Fish Diseases, 2(4):349–351, July 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1979:BRe

[Ano79e] Anonymous. Book reviews. Journal of Fish Diseases, 2(5):449– 453, September 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1979:E

[Ano79f] Anonymous. Errata. Journal of Fish Diseases, 2(3):262, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anderson:1979:CIR

[ARD79] D. P. Anderson, B. S. Robertson, and O. W. Dixon. Cellular immune response in rainbow trout Salmo gairdneri Richardson to Yersinia ruckeri O-antigen monitored by the passive haemolytic plaque assay test. Journal of Fish Diseases, 2(3):169–178, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Armas:1979:ODP

[Arm79] Gina Armas. Observations on diseases and parasites of mullet alevins Mugil cephalus L. from the Rio Moche coastal lagoon of Peru. Journal of Fish Diseases, 2(6):543–547, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bullock:1979:LAP

[BBN79] A. M. Bullock, R. S. Batty, and D. A. Neave. Lipophoromatosis in adult plaice Pleuronectes platessa L. Journal of Fish Diseases, 2(6):567–568, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bishop:1979:MBC

[BC79] R. K. Bishop and L. R. G. Cannon. Morbid behaviour of the commercial sand crab, Portunus pelagicus (L.), parasitized by Sacculina granifera Boschma, 1973 (Cirripedia: Rhizocephala). Journal of Fish Diseases, 2(2):131–144, March 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bach:1978:CSC

[BCC78] R. Bach, P. K. Chen, and G. B. Chapman. Changes in the spleen of the channel catfish Ictalurus punetatus Rafinesque indueed by infection with Aeromonas hydrophila. Journal of Fish Diseases, 1 (3):205–217, July 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bucke:1979:FSE

[BCC^{+}79] D. Bucke, G. D. Cawley, J. F. Craig, A. D. Pickering, and L. G. Willoughby. Further studies of an epizootic of perch Perca fluviatilis L., of uncertain aetiology. Journal of Fish Diseases, 2(4): 297–311, July 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bucke:1979:IPN

[BFMS79] D. Bucke, J. Finlay, D. Mcgregor, and C. Seagrave. Infectious pancreatic necrosis (IPN) virus: its occurrence in captive and wild fish in England and Wales. Journal of Fish Diseases, 2(6):549– 553, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Blasiola:1979:GHS

[Bla79] G. C. Blasiola, Jr. Glugea heraldi n. sp. (Microsporida, Glugeidae) from the seahorse Hippocampus erectus Perry. Journal of Fish Diseases, 2(6):493–500, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buchanan:1978:SHS

[BM78] J. S. Buchanan and C. R. Madeley. Studies on Herpesvirus scophthalmi infection of turbot Scophthalmus maximus (L.) ultrastructural observations. Journal of Fish Diseases, 1(4):283–295, October 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Becker:1979:HMF

[BO79] C. D. Becker and R. M. Overstreet. Haematozoa of marine fishes from the northern Gulf of Mexico. Journal of Fish Diseases, 2 (6):469–479, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bullock:1979:IUL

[BR79] A. M. Bullock and R. J. Roberts. Induction of UDN-like lesions in salmonid fish by exposure to ultraviolet light in the presence of phototoxic agents. Journal of Fish Diseases, 2(5):439–441, September 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bejerano:1979:MMS

[BSHR79] Y. Bejerano, S. Sarig, M. T. Horne, and R. J. Roberts. Mass mortalities in silver carp Hypophthalmichthys molitrix (Valenciennes) associated with bacterial infection following handling. Journal of Fish Diseases, 2(1):49–56, January 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bullock:1978:ERB

[BSS78] G. L. Bullock, H. M. Stuckey, and E. B. Shotts, Jr. Enteric redmouth bacterium: comparison of isolates from different geographic areas. Journal of Fish Diseases, 1(4):351–356, October 1978. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Blasiola:1979:AIS

[BT79] G. C. Blasiola, Jr. and J. C. Turnier. Algal infection of the sevengill shark, Notorynchus maculatus Ayres. Journal of Fish Diseases, 2(2):161–163, March 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buchanan:1978:CSN

[Buc78] J. S. Buchanan. Cytological studies on a new species of rickettsia found in association with a phage in the digestive gland of the marine bivalve mollusc, Tellina tenuis (da Costa). Journal of Fish Diseases, 1(1):27–43, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Blazer:1979:ELF

[BW79] V. S. Blazer and R. E. Wolke. An Exophiala-like fungus as the cause of a systemic mycosis of marine fish. Journal of Fish Diseases, 2(2):145–152, March 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheung:1979:SCM

[CNR79] P. J. Cheung, R. F. Nigrelli, and G. D. Ruggieri. Studies on cryptocaryoniasis in marine fish: effect of temperature and salinity on the reproductive cycle of Cryptocaryon irritans Brown, 1951. Journal of Fish Diseases, 2(2):93–97, March 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Conroy:1979:CGT

[CS79] D. A. Conroy and J. A. Santacana. A case of goitre and its treatment in a goldfish Carassius auratus (L.). Journal of Fish Diseases, 2(6):555–556, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carlisle:1979:IHN

[CSE79] J. C. Carlisle, K. A. Schat, and R. Elston. Infectious haematopoietic necrosis in rainbow trout Salmo gairdneri Richardson in a semi-closed system. Journal of Fish Diseases, 2(6):511–517, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Couch:1979:VDY

[CWHG79] J. A. Couch, J. T. Winstead, D. J. Hansen, and L. R. Goodman. Vertebral dysplasia in young fish exposed to the herbicide trifluralin. Journal of Fish Diseases, 2(1):35–42, January 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dorson:1978:IPN

[DCT78] M. Dorson, J. Castric, and C. Torchy. Infectious pancreatic necrosis virus of salmonids: biological and antigenic features of a pathogenic strain and of a non-pathogenic variant selected in RTG-2 cells. Journal of Fish Diseases, 1(4):309–320, October 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dykova:1978:TRT

[DL78] Iva Dyková and Jiří Lom. Tissue reaction of the three-spined stickleback Gasterosteus aculeatus L. to infection with Glugea anomala (Moniez, 1887). Journal of Fish Diseases, 1(1):83–90, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dykova:1979:HCT

[DL79] Iva Dyková and J. Lom. Histopathological changes in Trypanosoma danilewskyi Laveran & Mesnil, 1904 and Trypanoplasma borelli Laveran & Mesnil, 1902 infections of goldfish, Carassiw auratus (L.). Journal of Fish Diseases, 2(5):381–390, September 1979. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Doeson:1979:CDN

[DT79] M. Doeson and Corinne Torchy. Complement dependent neutralization of Egtved virus by trout antibodies. Journal of Fish Diseases, 2(4):345–347, July 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Duncan:1978:EOS

[Dun78] I. B. Duncan. Evidence for an oncovirus in swimbladder fibrosarcoma of Atlantic salmon Salmo salar L. Journal of Fish Diseases, 1(1):127–131, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Egidius:1978:HSP

[EA78] E. C. Egidius and K. Andersen. Host-specific pathogenicity of strains of Vibrio anguillarum isolated from rainbow trout Salmo gairdneri Richardson and saithe Pollachius virens (L.). Journal of Fish Diseases, 1(1):45–50, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Egidius:1979:UFA

[EA79a] E. Egidius and K. Andersen. The use of Furanace against vibriosis in rainbow trout Salmo gairdneri Richardson in salt water. Journal of Fish Diseases, 2(1):79–80, January 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Egidius:1979:BIP

[EA79b] E. C. Egidius and K. Andersen. Bath-immunization — a practical and non-stressing method of vaccinating sea farmed rainbow trout Salmo gairdneri Richardson against vibriosis. Journal of Fish Diseases, 2(5):405–410, September 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Edwards:1978:AIF

[Edw78] Clayton J. Edwards. Algal infections of fish tissue: a recent record and review. Journal of Fish Diseases, 1(2):175–179, April 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Emerson:1979:EAL

[ENB79] M. Emerson, B. L. Nicholson, and R. Bayer. Effects of Acholeplasma laidlawii and an unidentified mycoplasma on selected fish cell cultures and the replication of fish viruses. Journal of Fish Diseases, 2(3):227–238, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Eddy:1979:ICM

[ESB79] F. B. Eddy, G. R. Smart, and R. N. Bath. Ionic content of muscle and urine in rainbow trout Salmo gairdneri Richardson kept in water of high CO_{2} content. Journal of Fish Diseases, 2(2):105–110, March 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellis:1978:CAS

[EW78] A. E. Ellis and R. Wootten. Costiasis of Atlantic salmon, Salmo salar L. smelts in seawater. Journal of Fish Diseases, 1(4):389– 393, October 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:1979:EAP

[FB79] H. W. Ferguson and H. J. Ball. Epidemiological aspects of proliferative kidney disease amongst rainbow trout Salmo gairdneri Richardson in Northern Ireland. Journal of Fish Diseases, 2(3): 219–225, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:1979:STE

[Fer79] H. W. Ferguson. Scanning and transmission electron microscopical observations on Hexamita salmonis (Moore, 1922) related to mortalities in rainbow trout fry Salmo gairdneri Richardson. Journal of Fish Diseases, 2(1):57–67, January 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fijan:1978:IPN

[FG78] N. N. Fijan and G. Giorgetti. Infectious pancreatic necrosis: isolation of virus from eyed eggs of rainbow trout Salmo gairdneri Richardson. Journal of Fish Diseases, 1(3):269–270, July 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:1978:HFB

[FM78] H. W. Ferguson and D. H. Mccarthy. Histopathology of furunculosis in brown trout Salmo trutta L. Journal of Fish Diseases, 1(2): 165–174, April 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:1978:PKD

[FN78] H. W. Ferguson and E. A. Needham. Proliferative kidney disease in rainbow trout Salmo gairdneri Richardson. Journal of Fish Diseases, 1(1):91–108, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Foster:1978:FSP

[FSH78] C. A. Foster, T. G. Sarphie, and W. E. Hawkins. Fine structure of the peritrichous ectocommensal Zoothamnium sp. with emphasis on its mode of attachment to penaeid shrimp. Journal of Fish Diseases, 1(4):321–335, October 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gilbert:1979:VMT

[GGB79] J. P. Gilbert, J. B. Gratzek, and J. Brown. An in vitro method for testing synergistic action of parasiticides using malachite green and formalin as a model system. Journal of Fish Diseases, 2(3): 191–196, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hayunga:1979:OIP

[Hay79] E. G. Hayunga. Observations on the intestinal pathology caused by three caryophyllid tapeworms of the white sucker Catostomus commersoni Laécpède. Journal of Fish Diseases, 2(3):239–248, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huizinga:1979:HRS

[HEH79] H. W. Huizinga, G. W. Esch, and T. C. Hazen. Histopathology of red-sore disease (Aeromonas hydrophila) in naturally and experimentally infected largemouth bass Micropterus salmoides (Lacépède). Journal of Fish Diseases, 2(4):263–277, July 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hbtrick:1979:EWT

[HFK79] F. M. Hbtrick, J. L. Fryer, and M. D. Knittel. Effect of water temperature on the infection of rainbow trout Salmo gairdneri Richardson with infectious haematopoietic necrosis virus. Journal of Fish Diseases, 2(3):253–257, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Harbell:1979:SPV

[HHS79] S. C. Harbell, H. O. Hodgins, and M. H. Schiewe. Studies on the pathogenesis of vibriosis in Coho salmon Oncorhynchus kisutch (Walbaum). Journal of Fish Diseases, 2(5):391–404, September 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoffman:1979:CHK

[HKMS79] Glenn L. Hoffman, S. L. Kazubski, Andrew J. Mitchell, and C. E. Smith. Chilodonella hexasticha (Kiernik, 1909) (Protozoa, Ciliata) from North American warmwater fish. Journal of Fish Diseases, 2(2):153–157, March 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hedrick:1978:PIS

[HLF78] R. P. Hedrick, J. C. Leong, and J. L. Fryer. Persistent infections in salmonid fish cells with infectious pancreatic necrosis virus (IPNV). Journal of Fish Diseases, 1(4):297–308, October 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Harrison:1979:PHN

[HR79] J. G. Harrison and R. H. Richards. The pathology and histopathology of nephrocalcinosis in rainbow trout Salmo gairdneri Richardson in fresh water. Journal of Fish Diseases, 2(1):1–12, January 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

HaaStein:1978:MMA

[HSR78] T. HåStein, S. Jakob Saltveit, and R. J. Roberts. Mass mortality among minnows Phoxinus phoxinus (L.) in Lake Tveitevatn, Norway, due to an aberrant strain of Aeromonas salmonicida. Journal of Fish Diseases, 1(3):241–249, July 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoffman:1978:NMD

[HW78] G. L. Hoffman and R. Walker. A new myxosporidan, Dicauda atherinodi n.g., n. sp. of the minnow, Notropis atherinoides Rafinesque. Journal of Fish Diseases, 1(4):343–349, October 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hard:1979:HSF

[HWL79] G. C. Hard, R. Williams, and J. Lee. Histopathology of superficial fish ‘tumours’ found during a cancer survey of demersal species in Port Philip Bay, Victoria, Australia. Journal of Fish Diseases, 2 (6):455–467, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Inebi:1979:CFM

[IH79] W. Inebi and M. T. Horne. A complement-fixation method for the detection of Vibrio-specific antibodies in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 2(2):159–160, March 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kirmse:1978:HSS

[Kir78] P. D. Kirmse. Haemogregavina sachai n.sp. from cultured turbot Scophthabnus maximus (L.) in Scotland. Journal of Fish Diseases, 1(4):337–342, October 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kusuda:1978:SPS

[KK78] Riichi Kusuda and Hajime Kimura. Studies on the pathogenesis of streptococcal infection in cultured yellowtails Seriola spp.: the fate of Streptococcus sp. bacteria after inoculation. Journal of Fish Diseases, 1(1):109–114, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kimura:1979:SPS

[KK79] H. Kimura and R. Kusuda. Studies on the pathogenesis of streptococcal infection in cultured yellowtails Seriola spp.: effect of the cell free culture on experimental streptococcal infection. Journal of Fish Diseases, 2(6):501–510, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Katae:1979:DEB

[KKS^{+}79] H. Katae, K. Kouno, Y. Sekine, M. Hashimoto, and M. Shimizu. Distribution, excretion and biotransformation of ^{14}C-piromidic acid in goldfish Carassius aumtus (L.). Journal of Fish Diseases, 2(6):529–542, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Katae:1979:EPA

[KKT^{+}79] H. Katae, K. Kouno, Y. Takase, H. Miyazaki, M. Hashimoto, and M. Shimizu. The evaluation of piromidic acid as an antibiotic in fish: an in vitro and in vivo study. Journal of Fish Diseases, 2 (4):321–335, July 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lounatmaa:1978:EMU

[LJ78] K. Lounatmaa and J. Janatuinen. Electron microscopy of an ulcerative dermal necrosis (UDN)-like salmon disease in Finland. Journal of Fish Diseases, 1(4):369–375, October 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leatherland:1979:EDM

[LS79] J. F. Leatherland and R. A. Sonstegard. Effect of dietary mirex and PCB (Aroclor 1254) on thyroid activity and lipid reserves in rainbow trout Salmo gairdneri Richardson. Journal of Fish Diseases, 2(1):43–48, January 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mahajan:1979:EDI

[MAJK79] C. L. Mahajan, N. K. Agrawal, M. J. John, and V. P. Katta. Effect of the digenean Isoparorchis hypselobagri (Billet, 1898) on an airbreathing fish Channa punctatus (Bloch) with particular reference to biochemical and haematological changes. Journal of Fish Diseases, 2(6):519–528, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mcarthur:1978:HAP

[Mca78] C. P. Mcarthur. Humoral antibody production by New Zealand eels, against the intestinal trematode Telogaster opisthorchis Macfarlane, 1945. Journal of Fish Diseases, 1(4):377–387, October 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meyers:1979:PSC

[Mey79] T. R. Meyers. Preliminary studies on a chlamydial agent in the digestive diverticular epithelium of hard clams Mercenaria mercenaria (L.) from Great South Bay, New York. Journal of Fish Diseases, 2(3):179–189, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mccain:1979:TMD

[MGMW79] B. B. Mccain, W. D. Gronlund, M. S. Myers, and S. R. Wellings. Tumours and microbial diseases of marine fishes in Alaskan waters. Journal of Fish Diseases, 2(2):111–130, March 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mitchell:1978:NFP

[MGP78] A. J. Mitchell, J. M. Grizzle, and J. A. Plumb. Nifurpirinol (Furanace; P-7138) related lesions on channel catfish Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 1(1):115–121, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mills:1979:AFA

[Mil79] C. A. Mills. Attachment and feeding of the adult ectoparasitic digenean Transversotrema patialense (Soparkar, 1924) on the zebra fish Brachydanio rerio (Hamilton-Buchanan). Journal of Fish Diseases, 2(5):443–447, September 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:1978:ILG

[MR78] Kálmán Molnár and Miklós Reinhardt. Intestinal lesions in grasscarp Ctenopharyngodon idella (Valenciennes) infected with Balantidium ctenopharyngodonis Chen. Journal of Fish Diseases, 1(2): 151–156, April 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mcvicar:1979:FSN

[MW79] A. H. Mcvicar and P. G. White. Fin and skin necrosis of cultivated Dover sole Solea solea (L.). Journal of Fish Diseases, 2(6):557–562, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nicholson:1979:SHR

[NTJH79] B. L. Nicholson, G. W. Thorne, Caryle Janicki, and Anne Hanson. Studies on a host range variant from different isolates of infectious pancreatic necrosis virus (IPNV). Journal of Fish Diseases, 2(5):367–379, September 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

OLeary:1978:ETA

[OCF78] P. J. O’Leary, J. O. Cisar, and J. L. Fryer. The effect of temperature on agglutination activity of Coho salmon Oncorhynchus kisutch (Walbaum) antiserum. Journal of Fish Diseases, 1(1):123– 125, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Piavis:1979:TLO

[PBH79] G. W. Piavis, T. Beckerman, and J. H. Howell. Teratoid lesions and other developmental anomalies in the sea lamprey, Petromyzon marinus L. Journal of Fish Diseases, 2(5):353–366, September 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Phillips:1978:EOC

[PC78] W. J. Phillips and L. R. G. Cannon. Ecological observations on the commercial sand crab, Portunus pelagicus (L.), and its parasite, Sacculina granifera Boschma, 1973 (Cirripedia: Rhizocephala). Journal of Fish Diseases, 1(2):137–149, April 1978. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Peters:1979:ISG

[PP79] Gabriele Peters and Nicolaus Peters. The influence of salinity on growth and structure of epidermal papillomas of the European eel Anguilla anguilla L. Journal of Fish Diseases, 2(1):13–26, January 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Peters:1978:DST

[PPS^{+}78] Nicolaus Peters, Gabriele Peters, Hans F. Stich, Alf B. Acton, and Gisela Bresching. On differences in skin tumours of Pacific and Atlantic Flatfish. Journal of Fish Diseases, 1(1):3–25, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paperna:1978:USE

[PSC78] I. Paperna, I. Sabnai, and M. Castel. Ultrastructural study of epitheliocystis organisms from gill epithelium of the fish Sparus aurata (L.) and Liza ramada (Risso) and their relation to the host cell. Journal of Fish Diseases, 1(2):181–189, April 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Richards:1978:SHS

[RB78] R. H. Richards and J. S. Buchanan. Studies on Herpesvirus scophthalmi infection of turbot Scophthalmus maximus (L.): Histopathological observations. Journal of Fish Diseases, 1(3): 251–258, July 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:1979:PMC

[RB79] R. J. Roberts and A. M. Bullock. Papillomatosis in marine cultured rainbow trout Salmo gairdneri Richardson. Journal of Fish Diseases, 2(1):75–77, January 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Reynolds:1978:FRG

[RCC78] W. W. Reynolds, J. B. Covert, and Martha E. Casterlin. Febrile responses of goldfish Carassius auratus (L.) to Aeromonas hydrophila and to Escherichia coli endotoxin. Journal of Fish Diseases, 1(3): 271–273, July 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rau:1978:FDT

[RG78] M. E. Rau and D. M. Gordon. The frequency distribution of tetracotyles of Apatemon gracilis pellucidus (Yamaguti, 1933) in stickleback Culaea inconstans (Kirtland) populations of homogeneous age and size structure. Journal of Fish Diseases, 1(3):259–263, July 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rau:1979:BAD

[RGC79] M. E. Rau, D. M. Gordon, and M. A. Curtis. Bilateral asymmetry of Diplostomum infections in the eyes of lake whitefish Coregonm clupeaformis (Mitchill) and a computer simulation of the observed metacercarial distribution. Journal of Fish Diseases, 2(4):291–297, July 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Robekts:1978:BMF

[RH78] R. J. Robekts and M. T. Horne. Bacterial meningitis in farmed rainbow trout Salmo gairdneri Richardson affected with chronic pancreatic necrosis. Journal of Fish Diseases, 1(2):157–164, April 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roald:1979:LAS

[RH79] S. O. Roald and T. Håstein. Lymphosarcoma in an Atlantic salmon Salmo salar L. Journal of Fish Diseases, 2(3):249–251, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Richards:1978:ESI

[RHH78] R. H. Richards, A. Holliman, and S. Helgason. Exophiala salmonis infection in Atlantic salmon Salmo salar L. Journal of Fish Diseases, 1(4):357–368, October 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Robertson:1979:HPI

[Rob79] D. A. Robertson. Host-parasite interactions between Ichtyobodo necator (Henneguy, 1883) and farmed salmonids. Journal of Fish Diseases, 2(6):481–491, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Richards:1978:FDP

[RP78] R. H. Richards and A. D. Pickering. Frequency and distribution patterns of Saprolegnia infection in wild and hatchery-reared brown trout Salmo trutta L. and char Salvelinus alpinus (L.). Journal of Fish Diseases, 1(1):69–82, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Richards:1979:CSP

[RP79] R. H. Richards and A. D. Pickering. Changes in serum parameters of Saprolegnia-infected brown trout, Salmo trutta L. Journal of Fish Diseases, 2(3):197–206, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:1979:PRT

[RRB79] R. J. Roberts, R. H. Richards, and A. M. Bullock. Pansteatitis in rainbow trout Salmo gairdneri Richardson: a clinical and histopathological study. Journal of Fish Diseases, 2(2):85–92, March 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Scott:1979:PCF

[SG79] Alan L. Scott and John M. Grizzle. Pathology of cyprinid fishes caused by Bothriocephalus gowkongensis Yea, 1955 (Cestoda: Pseudophyllidea). Journal of Fish Diseases, 2(1):69–73, January 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smart:1979:NRT

[SKH^{+}79] G. R. Smart, D. Knox, J. G. Harrison, J. A. Ralph, R. H. Richard, and C. B. Cowey. Nephrocalcinosis in rainbow trout Salmo gairdneri Richardson; the effect of exposure to elevated CO_{2} concentrations. Journal of Fish Diseases, 2(4):279–289, July 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:1979:PLL

[Smi79] C. E. Smith. The prevention of liver lipoid degeneration (ceroidosis) and microcytic anaemia in rainbow trout Salmo gairdneri Richardson fed rancid diets: a preliminary report. Journal of Fish Diseases, 2(5):429–437, September 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Snieszko:1978:F

[Sni78] S. F. Snieszko. Foreword. Journal of Fish Diseases, 1(1):1, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sommerville:1978:HSB

[Som78] Christina Sommerville. The histopathology of Stephanochasmus baccatus Nicoll, 1907 in the digestive gland of Buccinum undatum (L.). Journal of Fish Diseases, 1(3):219–232, July 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:1979:HAT

[SPKM79] C. E. Smith, T. H. Peck, R. J. Klauda, and J. B. Mclaren. Hepatomas in Atlantic tomcod Microgadus tomcod (Walbaum) collected in the Hudson River estuary in New York. Journal of Fish Diseases, 2(4):313–319, July 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valtonen:1979:NRM

[Val79] E. Tellervo Valtonen. Neoechinorhynchus rutili (Müller, 1780) (Acanthocephala) in the whitefish Coregonm nasus (Pallas) sensu Svärdson from the Bay of Bothnia. Journal of Fish Diseases, 2(2): 99–103, March 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wales:1979:IHR

[Wal79] J. H. Wales. Induction of hepatoma in rainbow trout Salmo gaivdneri Richardson by the egg bath technique. Journal of Fish Diseases, 2(6):563–566, November 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Willoughby:1978:SSF

[Wil78] L. G. Willoughby. Saprolegnias of salmonid fish in Windermere: a critical analysis. Journal of Fish Diseases, 1(1):51–67, January 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wolf:1979:LVI

[WQCO79] K. E. Wolf, M. C. Quimby, C. V. Carlson, and W. J. Owens. Lymphocystis virus: infectivity of lesion preparations stored after lyophilization or simple desiccation. Journal of Fish Diseases, 2 (3):259–260, May 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wyatt:1978:SEM

[Wya78] E. J. Wyatt. Studies on the epizootiology of Myxobolus insidiosus Wyatt and Pratt, 1963 (Protozoa: Myxosporida). Journal of Fish Diseases, 1(3):233–240, July 1978. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yamamoto:1979:LEM

[YS79] T. Yamamoto and J. E. Sanders. Light and electron microscopic observations of sporogenesis in the myxosporida, Ceratomyxa shasta (Noble, 1950). Journal of Fish Diseases, 2(5):411– 428, September 1979. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).